OpenStudy (anonymous):

is the asymptotes of the following function and the curve lies about them.. f(x)=xsquare2/x-3

OpenStudy (ash2326):

@Muskan Is this your question? \[y=\frac{x^2}{x-3}\]

OpenStudy (anonymous):

yess

OpenStudy (ash2326):

I think the initial part of the question is missing, Could you post it?

OpenStudy (anonymous):

no is it

OpenStudy (anonymous):

onlyy

OpenStudy (ash2326):

ok, could you tell for what value of x the denominator is zero?

OpenStudy (anonymous):

i thnk (x-3)^2= (x-3)(x-3) X=3 is it rite??

OpenStudy (ash2326):

Is there (x-3)^2 in the denominator?

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

yeah x=3 is an asymptote

OpenStudy (ash2326):

Do you need to plot it also?

OpenStudy (anonymous):

thnxx

OpenStudy (experimentx):

\[ {x^2 \over (x - 3)^2} = 1+ {(x - 3)^2 - x^2 \over (x-3)^2}\] y = 1 is your horizontal asymptote